OpenStudy (anonymous):

SOMEONE HELP ME PLS MEDAL AND FAN Find the coordinates of the midpoint of if J (-9, 5) and K(21, -7). Find the slope of a line that has an x-intercept of -2 and a y-intercept of 6. Find the value of x so that the line passing through (x, 10) and (-4, 8) has a slope of 2/3

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

this is three different questions right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

Find the coordinates of the midpoint of if J (-9, 5) and K(21, -7). take the average in each coordinate, i.e \[\left(\frac{21-9}{2},\frac{5-7}{2}\right)\]

OpenStudy (anonymous):

alright

OpenStudy (misty1212):

compute and get \((6,-1)\)

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

Find the slope of a line that has an x-intercept of -2 and a y-intercept of 6. the two points are \((-2,0)\) and \((0,6)\)

OpenStudy (misty1212):

slope is \[m=\frac{6-0}{0-(-2)}=\frac{6}{2}\]

OpenStudy (misty1212):

better known as 3

OpenStudy (misty1212):

Find the value of x so that the line passing through (x, 10) and (-4, 8) has a slope of 2/3 put \[\frac{10-8}{x+4}=\frac{2}{3}\] and solve for \(x\)

OpenStudy (misty1212):

you get \[\frac{2}{x+4}=\frac{2}{3}\]so \[x+4=3\] solve that one

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

x = -1?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i have to 2 more

OpenStudy (misty1212):

ohh goodie!

OpenStudy (anonymous):

Write the equation of a line in slope intercept form that is parallel to the line y= 1/3x+5 and passes through (-9, 5).

OpenStudy (misty1212):

on line class?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ick

OpenStudy (misty1212):

parallel line also has slope \(\frac{1}{3}\)

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form from the graph below.

OpenStudy (misty1212):

wait we are not done with the first one yet!!

OpenStudy (misty1212):

Write the equation of a line in slope intercept form that is parallel to the line y= 1/3x+5 and passes through (-9, 5). slope is \(\frac{1}{3}\) point is \((-9,5)\) equation is \[y-5=\frac{1}{3}(x+9)\] you can multiply out and then solve for \(y\)

OpenStudy (misty1212):

for the picture you need the slope and the y intercept do you see the y intercept?

OpenStudy (anonymous):

2?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

how about the slope? do you see that?

OpenStudy (anonymous):

3/1 ?

OpenStudy (misty1212):

on no

OpenStudy (misty1212):

one step to the right, four steps up slope is \(\frac{4}{1}=4\)

OpenStudy (misty1212):

so equation is \[y=4x+2\]

OpenStudy (anonymous):

oh ok